OpenStudy (anonymous):

Jacqueline earned a twelve percent pay raise. If she was previously making 9 dollars and fifty cents per hour, how much will she make per hour now?

OpenStudy (anonymous):

A = 9.50 r = 12.5 \[Final Amount = A \times \left(1 + \frac{ r }{ 100 } \right)\]